23g ethanol is dissolved in 36g water . Find mole fraction of ethanol ?

Grade:12th pass

3 Answers

Arun
6 years ago
Dear Deepika
 
Ethanol = C2H5OH
molecular mass = 24 + 5 + 16 +5 = 50
 
moles of Ethanol = 23/50
 
moles of water = 36/18 = 2 moles
 
now mole fraction of ethanol = (23/50) / (23/50 + 2)
 = 23/123
 
 
Regards
Arun (askIITians forum expert)
Vibhuti
5 years ago
Molar mass of ethanol= 50
Moles of ethanol=23/50
Moles of water= 36/18 =2 
Now mole fraction of ethanol= (23/50)/(23/50+2)
                                                     =23/123
 
Anjali kumari
5 years ago
As mile=given mass/ molecular mass.Therefore the mile of ethanol is 23/50 moles. And the mole of water =2 moles. So mole fraction of ethanol=(23/50)/(23/50+2).